# Hallucination Detection in Foundation Models for Decision-Making: A Flexible Definition and Review of the State of the Art

### v1

@@ 3.3 Why do they happen?
There are several speculations as to how hallucinations come about during deployment. First and foremost, like any learning task, foundation models are sensitive to biases in training data Rawte et al. (2023). Once a model is trained on a given large dataset, some facts may become out-of-date or stale at any point in time Puthumanaillam et al. (2024). Furthermore, as the training set is embedded into a smaller encoding dimension, the knowledge within an L(V)LM’s frozen parameters is lossy, and models cannot feasibly be fine-tuned every time there is new data Peng et al. (2023); Elaraby et al. (2023). Zhang et al. (2023b) recommend changing algorithm parameters at runtime, such as, temperature (spread of probability distribution of next token), top-$K$ sampling (narrows the set of next tokens to be considered), and beam search (choosing a set of possible beams, i.e., trajectories, of next tokens based on high conditional probabilities), but the process of tuning these parameters is expensive.
@@
To combat out-of-date training data, some works provide models with an external knowledge-base of information to pull facts from, with the hope of increasing model accuracy. Even with this up-to-date information, Zhang et al. (2023c) pose that there may exist a misalignment between the true capabilities of a model, and what a user believes the model is capable of, leading to poor prompt engineering. In fact, poor prompting is one of the most significant causes of hallucinations. Chen et al. (2021) find that poor quality prompts lead to poor quality generations, in the context of code completion. This phenomenon is attributed to the reconstruction training objective of LLMs attempting to maximize the likelihood of next generated tokens, given context and past outputs, i.e.,
@@
While foundation models condition generated tokens on ground-truth text without hallucinations at train time, during inference, the model chooses future tokens conditioned on previously (possibly hallucinated) generated text. As such, Chen et al. (2023d) and Varshney et al. (2023) state that generated outputs are more likely to contain hallucinations if prior tokens are hallucinated as well. Furthermore, Li et al. (2023a) find that, even if prompt context provided to a foundation model is relevant, the model may choose to ignore the information and revert to its own (possibly outdated or biased) parameterized knowledge.
@@
4 Detection and Mitigation Strategies
Table 2: A summary of hallucination detection & mitigation methods discussed in Section 4. Deployment scenarios are split into question-answering (QA), information retrieval (IR), image captioning (IC), image generation (IG), & planning (P) tasks. The method ID includes the subsection the method appears in the paper and the order in which it appears in the subsection. Bolded method IDs are deployed to decision-making tasks specifically. Custom datasets, custom simulators, & real-world experiments for testing are abbreviated as CD, CS, & RW, respectively.
Modality Method Type Method ID Application Deployment Evaluation Datasets or Simulators
Detection Mitigation
White-box Hidden States 4.1.1.1 $\bullet$ QA CD
4.1.1.2 $\bullet$ QA CD
4.1.1.3 $\bullet$ QA DecodingTrust Wang et al. (2023a)
TruthfulQA Lin et al. (2022b)
Attention Weights 4.1.2.1 $\bullet$ $\bullet$ IC MSCOCO Lin et al. (2014)
Visual Genome Krishna et al. (2017)
Honesty Alignment 4.1.3.1 $\bullet$ QA CD
4.1.3.2 $\bullet$ $\bullet$ QA TriviaQA Joshi et al. (2017)
Grey-box Concept Probabilities 4.2.1.1 $\bullet$ $\bullet$ IR/QA HotpotQA Yang et al. (2018)
CD
4.2.1.2 $\bullet$ IC MSCOCO Lin et al. (2014)
4.2.1.3 $\bullet$ P Ravens Zeng et al. (2021)
BabyAI Chevalier-Boisvert et al. (2019)
VirtualHome Puig et al. (2018)
Conformal Prediction 4.2.2.1 $\bullet$ $\bullet$ IR/QA MIMIC-CXR Johnson et al. (2019)
CNN/DM Hermann et al. (2015)
TriviaQA Joshi et al. (2017)
4.2.2.2 $\bullet$ $\bullet$ QA MMLU Hendrycks et al. (2021)
4.2.2.3 $\bullet$ $\bullet$ P KnowNo TableSim Ren et al. (2023)
RW
4.2.2.4 $\bullet$ $\bullet$ P KnowNo TableSim Ren et al. (2023)
CD
4.2.2.5 $\bullet$ $\bullet$ P CS
Black-box Analyzing Samples 4.3.1.1 $\bullet$ IR/QA CD
4.3.1.2 $\bullet$ $\bullet$ IR/QA CD
4.3.1.3 $\bullet$ $\bullet$ IR/QA GSM8888Cobbe et al. (2021)
MMLU Hendrycks et al. (2021)
CD
4.3.1.4 $\bullet$ $\bullet$ P SaGC Park et al. (2024)
CS
RW
4.3.1.5 $\bullet$ $\bullet$ IR/QA CD
4.3.1.6 $\bullet$ $\bullet$ IR/QA Quest Malaviya et al. (2023)
MultiSpanQA Li et al. (2022)
FActScore Min et al. (2023)
CD
4.3.1.7 $\bullet$ IC/QA MSCOCO Lin et al. (2014)
A-OKVQA Schwenk et al. (2022)
GQA Hudson and Manning (2019)
4.3.1.8 $\bullet$ QA BIG-Bench Srivastava et al. (2023)
GSM8888Cobbe et al. (2021)
MMLU Hendrycks et al. (2021)
Adversarial Prompting 4.3.2.1 $\bullet$ IG CD
4.3.2.2 $\bullet$ $\bullet$ IR/QA Natural Questions Kwiatkowski et al. (2019)
CD
Proxy Model 4.3.3.1 $\bullet$ IR/QA CD
4.3.3.2 $\bullet$ IR/QA SQuAD Rajpurkar et al. (2016)
HotpotQA Yang et al. (2018)
TriviaQA Joshi et al. (2017)
4.3.3.3 $\bullet$ IR/QA CD
Grounding Knowledge 4.3.4.1 $\bullet$ IR/QA Natural Questions Kwiatkowski et al. (2019)
StrategyQA Geva et al. (2021)
QRreCC Anantha et al. (2021)
4.3.4.2 $\bullet$ $\bullet$ IR/QA LC-QuAD Trivedi et al. (2017)
KQA-Pro Cao et al. (2022)
ScienceQA Lu et al. (2022)
4.3.4.3 $\bullet$ $\bullet$ IR/QA FuzzyQA Zhang et al. (2023c)
4.3.4.4 $\bullet$ $\bullet$ IR/QA CD
4.3.4.5 $\bullet$ P TextWorld Côté et al. (2019)
Constraint Satisfaction 4.3.5.1 $\bullet$ $\bullet$ P CD
4.3.5.2 $\bullet$ P RoboEval Hu et al. (2024)
Hallucination detection and mitigation methods can be classified into three types (white-, grey-, and black-box) depending on the available inputs to the algorithm. Generally, given some context, a foundation model outputs a predicted sequence of tokens, the corresponding probabilities of each token, and embeddings of the generation from intermediate layers in the network. White-box hallucination detection methods assume access to all three output types, grey-box require token probabilities, and black-box only need the predicted sequence of tokens. Because not all foundation models provide access to their hidden states, or even the output probability distribution of tokens (e.g., the ChatGPT web interface), black-box algorithms are more flexible during testing. In this section, we present existing detection and mitigation approaches clustered by input type. While several of these works show promise in QA and object detection settings, many of them require further validation on decision-making tasks, and we will point out these methods as they come about. Works in this section are summarized in Table 2.
4.1 White-box Methods
@@ 4.1.1 Hidden States
    Azaria and Mitchell (2023) empirically find that language models attempt to correct themselves after outputting an untruthful claim. As such, they hypothesize that the internal states of a model must have some understanding of whether the output is correct. Furthermore, the authors stray away from directly using the token probabilities, even though they have correlation with model accuracy Varshney et al. (2023), because the complete output sentence’s probability is dependent on the length of the generation and appearance frequency of tokens. Azaria and Mitchell (2023) present SAPLMA, a simple classifier trained with supervised learning, that takes the activation values of a hidden layer of an LLM as input, and outputs the probability of the generated claim being true. SAPLMA is shown to be able to identify untruthful outputs, even when trained on a held-out dataset on a completely different topic than evaluated on.
  2. 2.

### v2

@@ 3.3. Why do they happen?
There are several speculations as to how hallucinations come about during deployment. First and foremost, like any learning task, foundation models are sensitive to biases in training data (Rawte et al., 2023). Once a model is trained on a given large dataset, some facts may become out-of-date or stale at any point in time (Puthumanaillam et al., 2024). Furthermore, as the training set is embedded into a smaller encoding dimension, the knowledge within an L(V)LM’s frozen parameters is lossy, and models cannot feasibly be fine-tuned every time there is new data (Peng et al., 2023; Elaraby et al., 2023). Zhang et al. (2024b) recommend changing algorithm parameters at runtime, such as, temperature (spread of probability distribution of next token), top-$K$ sampling (narrows the set of next tokens to be considered), and beam search (choosing a set of possible beams, i.e., trajectories, of next tokens based on high conditional probabilities), but the process of tuning these parameters is expensive.
@@
To combat out-of-date training data, some works provide models with an external knowledge-base of information to pull facts from, with the hope of increasing model accuracy. Even with this up-to-date information, Zhang et al. (2024a) pose that there may exist a misalignment between the true capabilities of a model, and what a user believes the model is capable of, leading to poor prompt engineering. In fact, poor prompting is one of the most significant causes of hallucinations. Chen et al. (2021) find that poor quality prompts lead to poor quality generations, in the context of code completion. This phenomenon is attributed to the reconstruction training objective of LLMs attempting to maximize the likelihood of next generated tokens, given context and past outputs (Mialon et al., 2023)i.e.,
@@
While foundation models condition generated tokens on ground-truth text without hallucinations at train time, during inference, the model chooses future tokens conditioned on previously (possibly hallucinated) generated text. As such, Chen et al. (2023a) and Varshney et al. (2023) state that generated outputs are more likely to contain hallucinations if prior tokens are hallucinated as well. Furthermore, Li et al. (2023c) find that, even if prompt context provided to a foundation model is relevant, the model may choose to ignore the information and revert to its own (possibly outdated or biased) parameterized knowledge.
@@
4. Detection and Mitigation Strategies
Hallucination detection and mitigation methods can be classified into three types (white-, grey-, and black-box) depending on the available inputs to the algorithm. Generally, given some context, a foundation model outputs a predicted sequence of tokens, the corresponding probabilities of each token, and embeddings of the generation from intermediate layers in the network. White-box hallucination detection methods assume access to all three output types, grey-box require token probabilities, and black-box only need the predicted sequence of tokens. Because not all foundation models provide access to their hidden states, or even the output probability distribution of tokens (e.g., the ChatGPT web interface), black-box algorithms are more flexible during testing. In this section, we present existing detection and mitigation approaches clustered by input type. While several of these works show promise in QA and image captioning settings, many of them require further validation on decision-making tasks, and we will point out these methods as they come about. Works in this section are summarized in Table LABEL:tab:summary. We also provide additional details about the evaluation setups of works deployed on custom datasets, simulators, or the real-world. Certain methods that are less related to decision-making are described in Appendix C. Additionally, frequently used metrics, datasets, and simulators are summarized in Appendix D.
4.1. White-box Methods
@@ 4.1.1. Hidden States
    Azaria and Mitchell (2023) empirically find that language models attempt to correct themselves after outputting an untruthful claim. As such, they hypothesize that the internal states of a model must have some understanding of whether the output is correct. Furthermore, the authors stray away from directly using the token probabilities, even though they have correlation with model accuracy (Varshney et al., 2023), because the complete output sentence’s probability is dependent on the length of the generation and appearance frequency of tokens. Azaria and Mitchell present SAPLMA, a simple classifier trained with supervised learning, that takes the activation values of a hidden layer of an LLM as input, and outputs the probability of the generated claim being true. The authors choose to collect a new dataset of simple statements with corresponding true/false answers, since popular datasets like FEVER (Thorne et al., 2018) do not partition statements by topic, and some statements cannot be cleanly classified. In total, their dataset contains $6$K sentences spanning six topics. SAPLMA is shown to be able to identify untruthful outputs, even when trained on a held-out dataset on a completely different topic than evaluated on.
  2. (2)